copy of email to Fred:
Many problems at OSC today. You guessed it Gen Test! When they transferred back they took out a breaker for the 120V transformer that serves OSC. Everything went down. Now as the Metasys devices come back up, we are discovering everything moving way to slow. All the OR reheat valves move very slow and one of the biggest problems of the morning was the steam valve for the heat exchanger. It would not move but a percent every five minutes, so we lost all the OR’s and they had to delay Surgeries. We got it all back up with overrides and monitoring.
Do you think Kyle’s changes got blown out? I also noticed that there was no reset schedule in, it was at 120* high and low. I put the “output Low at 140” to try and help the temps come up with a little warmer water. (I actually had it at 160* but we overshot a couple rooms, so I bumped back down).